# Approximating evidence via bounded harmonic means

### v1

@@ Abstract
Efficient Bayesian model selection relies on the model evidence or marginal likelihood, whose computation often requires evaluating an intractable integral. The harmonic mean estimator (HME) has long been a standard method of approximating the evidence. While computationally simple, the version introduced by Newton and Raftery [14] potentially suffers from infinite variance. To overcome this issue, Gelfand and Dey [4] defined a standardized representation of the estimator based on an instrumental function and Robert and Wraith [20] later proposed to use higher posterior density (HPD) indicators as instrumental functions. Following this approach, a practical method is proposed, based on an elliptical covering of the HPD region with non-overlapping ellipsoids. The resulting estimator (ECMLE) not only eliminates the infinite-variance issue of the original HME and allows exact volume computations, but is also able to be used in multimodal settings. Through several examples, we illustrate that ECMLE outperforms other recent methods such as THAMES and Mixture THAMES [11]. Moreover, ECMLE demonstrates lower variance—a key challenge that subsequent HME variants have sought to address—and provides more stable evidence approximations, even in challenging settings.
@@ 1 Introduction
for expanding HMEs (known as the Gelfand–Dey estimator) by allowing for more flexible instrumental density functions $\varphi(\theta)$ than the prior density $\pi(\theta)$. The Gelfand-Dey’s methodology further applies to complex models including hierarchical and non-conjugate models. However, the accuracy of their estimators heavily depends on the choice of the instrumental function $\varphi(\theta)$. As a result, Robert and Wraith [20] and Marin and Robert [8] later introduced HPD-truncated estimators based on the representation (1), which select the instrumental density as a uniform distribution over approximate highest posterior density (HPD) regions. The authors based their approach on the convex hull of MCMC simulations with the highest density values, ensuring boundedness of the importance ratio in (1) but requiring a computationally costly derivation of the convex hull. Similarly, Weinberg [22] used an indicator function as part of the instrumental function so as to limit the integration domain to a well-sampled region, thereby improving the stability and accuracy of the marginal likelihood estimators. However, Weinberg [22]’s approach becomes computationally intensive for complex posteriors such as heavy tail or multimodal cases. Wang et al. [21] proposed a weighted sum of indicator functions over a partition of the parameter space leading to more stable and consistent estimates of the evidence. The ensuing method is however more computationally intensive than a simple estimator since it requires a density estimation within each partition and adds costly post-processing. Furthermore, in the case of complex posteriors, the partitioning of the space may fail to capture accurately the structure of the posterior.
@@
Closer to our proposal, Reichl [17] constructed a geometrically motivated marginal likelihood estimator that leverages posterior draws and an ellipsoidal region centered on the posterior mode (MAP) to compute the evidence stably and efficiently, i.e., avoiding the high variance of classical estimators. The ellipsoid central to the method is derived from the empirical covariance of the posterior draws and most crucially, it enables an exact analytic calculation of its volume. More recently, Metodiev et al. [12] adopted a similar approach they called THAMES, where the ellipsoid is derived from a Gaussian distribution centered at the posterior mean or MAP, while within approximate $\alpha\%$-HPD regions. The covariance matrix behind the ellipsoid is estimated from an MCMC posterior sample and the method seeks to minimize the estimator’s variance by selecting an optimal ellipsoid radius to balance bias versus variance. Due to the simple and closed form of their estimators, Reichl [17]’s and Metodiev et al. [12]’s methods are practical and easy to implement. Their application is, however, limited to unimodal, smooth, and roughly Gaussian-shaped posteriors with low to moderate dimensions. Indeed, in multimodal or strongly skewed posterior cases, the ellipsoidal truncation may miss significant posterior regions and include irrelevant, low-posterior-density areas. To overcome these difficulties, Metodiev et al. [11] subsequently extended THAMES for multivariate mixture models, addressing key limitations of the earlier method, but the efficiency of their approach still depends on adaptation to the geometry of the posterior region and a potential computational cost of using auxiliary simulations to evaluate complex intersection volumes.
@@ 2 Approximating the evidence by harmonic means
and it holds for all probability density functions $\varphi(\cdot)$ that are well-defined over the posterior support. This identity justifies the following estimator
@@
2.0.1 Instrumental prior distribution
@@
2.0.2 Gaussian instrumental distribution
@@
$V(A)=\frac{\pi^{d/2}r^{d}|\hat{\Sigma}|^{-\frac{1}{2}}}{\Gamma(d/2+1)},$ (4)
@@
2.0.3 Uniform instrumental distribution for a convex hull of $\alpha$-HPD samples
In order to bound the variance of the estimator (2), Robert and Wraith [20] proposed choosing the instrumental function $\varphi(\cdot)$ as a uniform distribution over a HPD region of coverage $\alpha\%$ (called an $\alpha$-HPD region). This region is approximated from the $\alpha\%$ fraction of a simulated posterior sample that corresponds to the largest arguments of $\pi(\theta)L(\theta)$ by constructing a convex hull, $\mathfrak{H}_{\alpha}$. The associated estimator (2) is then
@@ 2.0.3 Uniform instrumental distribution for a convex hull of $\alpha$-HPD samples
The authors illustrated the efficiency of the method on a 22-dimensional toy example. However, the extension to higher dimensions is uncertain, due to the computational challenge of deriving the volume of the convex hull. In addition, the inclusion of the hull into an HPD region is not guaranteed and may be inefficient.
2.0.4 Uniform instrumental distribution over a Gaussian ellipsoid
Related to the proposal of DiCiccio et al. [3], Metodiev et al. [12] developed an instrumental distribution called Truncated Harmonic Mean Estimator (with the acronym THAMES). THAMES is a combination of both methods previously mentioned and uses a uniform distribution over the ellipsoid $A$ defined in (3). Since the volume of the ellipsoid $A$ is available as (4), THAMES bypasses both the computing issue of the volume and the derivation of the convex hull of Robert and Wraith [20]. Like the previous estimators, THAMES provides an unbiased estimator of $Z^{-1}$, provided that the posterior density does not vanish within the ellipsoid $A$. While the choice of the radius $r$ towards minimizing the variance of $\hat{Z}^{-1}$ is argued to be about $\sqrt{d+1}$ as in the Gaussian case, there is no guarantee that $A$ is included within a HPD region. Therefore, it faces limitations in multimodal cases (a case illustrated later on Figure 7), which may invalidate THAME’s consistency and reliability in these settings.
2.0.5 Uniform instrumental distribution over a truncated Gaussian ellipsoid
Most recently, Metodiev et al. [11] proposed a new version of THAMES, which we denote by Mix-THAMES for simplicity, and is specifically tailored for multimodal posterior distributions. This updated method relies on Reichl [17]’s proposal and includes a truncation strategy, towards using a support $A^{\prime}_{\hat{\theta},\hat{\Sigma},r,\alpha}$ made of the intersection of the ellipsoid $A$ in (3), and a $50\%$-HPD region. The THAMES for a mixture of $j=1,\ldots,J$ distributions of the same family,
@@ 2.0.5 Uniform instrumental distribution over a truncated Gaussian ellipsoid
$A^{\prime}_{\hat{\theta},\hat{\Sigma},r,\alpha}=\{\theta:(\theta-\hat{\theta})^{T}\hat{\Sigma}^{-1}(\theta-\hat{\theta})<r^{2},\pi(\theta)L(\theta)>\hat{q}_{\alpha}\}$
@@ 3 ECMLE method
The ECMLE estimator is based on the Gelfand-Dey identity (1) and the framework of Robert and Wraith [20] (5), where a uniform distribution over a convex hull is used as the instrumental function. In contrast, our method constructs the set $\mathcal{E}$ as a union of locally adapted ellipsoids, enabling exact volume computation while effectively capturing multimodal and irregularly shaped posteriors.
@@
Unbiased estimation requires that the region $\mathcal{E}$ be constructed independently of the sample used for evaluation. Given a collection of posterior draws $\{\boldsymbol{\theta}_{i}\}_{i=1}^{2T}$ from either MCMC or independent sampling, we partition them into two equal subsets. The first subset $\{\boldsymbol{\theta}_{t}\}_{t=1}^{T}$ defines the HPD region, while the second $\{\boldsymbol{\theta}_{t}\}_{t=T+1}^{2T}$ evaluates the estimator.
With respect to the chosen HPD level $\alpha$, we determine a threshold $c$ on the unnormalized posterior densities that separates the high- and low-density samples (Algorithm 1). This yields the sets of high-density points $\boldsymbol{\Theta}_{\mathrm{HPD}}=\{\boldsymbol{\theta}_{t}:\tilde{\pi}(\theta_{t}|{x})\geq c,\ t\leq T\}$ and low-density points $\boldsymbol{\Theta}_{\mathrm{LPD}}$. The threshold $c$ serves as the target boundary that ellipsoids must respect.
@@
1:Input: HPD samples $\boldsymbol{\Theta}_{HPD}$ with posterior values, LPD samples $\boldsymbol{\Theta}_{LPD}$, posterior function $\pi(\boldsymbol{\theta})$, HPD level $\alpha$
2:Output: Union of non-overlapping ellipsoids, $\mathcal{E}$, and total volume $V_{\text{total}}$
@@
6:Initialize: $\mathcal{E}\leftarrow\emptyset$, $V_{\text{total}}\leftarrow 0$, $\mathrm{Centers}\leftarrow$ subsampled $\boldsymbol{\Theta}_{HPD}$
@@
27:  $V_{\text{total}}\leftarrow V_{\text{total}}+v$
@@
31:Return $\mathcal{E},\,V_{\text{total}}$
We first subsample $\boldsymbol{\Theta}_{\mathrm{HPD}}$ by a factor $k$ (typically 0.05–0.1) to obtain candidate centers, ordered by posterior density to prioritize high-probability regions. For each candidate $\boldsymbol{\theta}^{*}$, we determine the ellipsoid shape through the following procedure:
The primary axis direction $\mathbf{u}_{1}$ points toward the nearest low-density point, capturing the dominant direction of posterior decay. Along this axis, we find radius $r_{1}$ where $\tilde{\pi}(\boldsymbol{\theta}^{*}+r_{1}\mathbf{u}_{1}|x)=c$ via a bisection search. An orthogonal basis $\{\mathbf{u}_{1},\ldots,\mathbf{u}_{d}\}$ is found using Gram-Schmidt, and semi-axes lengths $s_{i}$ are determined by finding the HPD boundary along each direction, yielding a shape matrix $\boldsymbol{\Sigma}=\mathbf{U}\,\mathrm{diag}(s_{1}^{2},\ldots,s_{d}^{2})\,\mathbf{U}^{\top}$.
To preserve the non-overlapping property required for volume computation, we adopt a conservative proximity rule: ellipsoids whose centers are closer than the sum of their effective radii are rejected. This guarantees disjointness and enables the exact computation of $V(\mathcal{E})=\sum_{j}\frac{\pi^{d/2}}{\Gamma(d/2+1)}\sqrt{\det(\boldsymbol{\Sigma}_{j})}$.
@@
With the ellipsoid collection $\mathcal{E}$ and total volume $V_{\text{total}}$ fixed, the marginal likelihood is evaluated using the second sample. For each parameter vector $\{\boldsymbol{\theta}_{t}\}_{t=T+1}^{2T}$, membership in $\mathcal{E}$ is determined by verifying whether it falls inside any ellipsoid $\mathfrak{e}_{j}$. This is assessed based on the Mahalanobis distance between $\boldsymbol{\theta}_{t}$ and the ellipsoid center $\boldsymbol{\mu}_{j}$, computed using the covariance matrix $\boldsymbol{\Sigma}_{j}$. The corresponding estimator is summarized in Algorithm 3.
@@
1:Input: Second sample $\{\boldsymbol{\theta}_{t}\}_{t=T+1}^{2T}$ with unnormalized posterior densities $\{\tilde{\pi}(\theta_{t}|{x})\}_{t=T+1}^{2T}$, ellipsoid collection $\mathcal{E}$, total volume $V_{total}$
@@
$\hat{Z}^{-1}=\frac{1}{T}\sum_{t=T+1}^{2T}\frac{\mathbf{1}_{\mathcal{E}}(\theta_{t})}{V_{\mathrm{total}}\;\tilde{\pi}(\theta_{t}\mid x)}.$
@@
The computational efficiency of ECMLE is critical for practical applications. We analyze the time complexity of each algorithm component, demonstrating that the method scales favorably with sample size and dimension.
@@
4 Illustrations
In this section, we empirically compare ECMLE with THAMES estimators through several examples. We focus on THAMES because it represents the current state of the art among harmonic-mean–based estimators. THAMES and its mixture variant (Mix-THAMES) were recently shown to outperform earlier bounded harmonic mean approaches and to provide stable, closed-form evidence approximations across a range of models. As ECMLE builds directly on the same harmonic-mean identity and HPD-based truncation principle, this comparison allows a fair and direct assessment within the same methodological family.
We use the notation $\boldsymbol{x}=\{x_{1},\ldots,x_{n}\}$ to indicate a set of $n$ observations. Each experiment was performed using 100 replications of Algorithms 1–3 for all examples, and the number of MCMC draws was set to $T=10^{5}$
4.0.1 Example 1: Multivariate Gaussian distributions

### v2

@@ Abstract
Efficient Bayesian model selection relies on the model evidence or marginal likelihood, whose computation often requires evaluating an intractable integral. The harmonic mean estimator (HME) has long been a standard method of approximating the evidence. While computationally simple, the version introduced by Newton and Raftery [15] potentially suffers from infinite variance. To overcome this issue, Gelfand and Dey [4] defined a standardized representation of the estimator based on an instrumental function and Robert and Wraith [21] later proposed to use higher posterior density (HPD) indicators as instrumental functions. Following this approach, a practical method is proposed, based on an elliptical covering of the HPD region with non-overlapping ellipsoids. The resulting estimator, called the Elliptical Covering Marginal Likelihood Estimator (ECMLE), not only eliminates the infinite-variance issue of the original HME and allows exact volume computations, but is also able to be used in multimodal settings. Through several examples, we illustrate that ECMLE outperforms other recent methods such as THAMES and its improved version [12]. Moreover, ECMLE demonstrates lower variance—a key challenge that subsequent HME variants have sought to address—and provides more stable evidence approximations, even in challenging settings.
@@
ripThis paper is dedicated to the memory of Karim Benabed, who collaborated with CPR and DW on evidence approximation in [8] and who died in a cycling accident in December 2025.
@@ 1 Introduction
for expanding HMEs (known as the Gelfand–Dey estimator) by allowing for more flexible instrumental density functions $\varphi(\theta)$ than the prior density $\pi(\theta)$. The Gelfand-Dey’s methodology further applies to complex models including hierarchical and non-conjugate models. However, the accuracy of their estimators heavily depends on the choice of the instrumental function $\varphi(\theta)$. As a result, Robert and Wraith [21] and Marin and Robert [9] later introduced HPD-truncated estimators based on the representation (1), which select the instrumental density as a uniform distribution over approximate highest posterior density (HPD) regions. The authors based their approach on the convex hull of MCMC simulations with the highest density values, ensuring boundedness of the importance ratio in (1) but requiring a computationally costly derivation of the convex hull. Similarly, Weinberg [23] used an indicator function as part of the instrumental function so as to limit the integration domain to a well-sampled region, thereby improving the stability and accuracy of the marginal likelihood estimators. However, Weinberg [23]’s approach becomes computationally intensive for complex posteriors such as heavy tail or multimodal cases. Wang et al. [22] proposed the Partition Weighted Kernel (PWK) estimator, based on a weighted sum of indicator functions over a partition of the parameter space leading to more stable and consistent estimates of the evidence. The ensuing method is however more computationally intensive than a simple estimator since it requires a density estimation within each partition and adds costly post-processing. Furthermore, in the case of complex posteriors, the partitioning of the space may fail to capture accurately the structure of the posterior.
@@
Closer to our proposal, Reichl [18] constructed a geometrically motivated marginal likelihood estimator that leverages posterior draws and an ellipsoidal region centered on the posterior mode (MAP) to compute the evidence stably and efficiently, i.e., avoiding the high variance of classical estimators. The ellipsoid central to the method is derived from the empirical covariance of the posterior draws and most crucially, it enables an exact analytic calculation of its volume. More recently, Metodiev et al. [13] adopted a similar approach they called the Truncated Harmonic Mean Estimator (THAMES), where the ellipsoid is derived from a Gaussian distribution centered at the posterior mean or MAP, while within approximate $\alpha$-HPD regions. The covariance matrix behind the ellipsoid is estimated from an MCMC posterior sample and the method seeks to minimize the estimator’s variance by selecting an optimal ellipsoid radius to balance bias versus variance. Due to the simple and closed form of their estimators, Reichl [18]’s and Metodiev et al. [13]’s methods are practical and easy to implement. Their application is, however, limited to unimodal, smooth, and roughly Gaussian-shaped posteriors with low to moderate dimensions. Indeed, in multimodal or strongly skewed posterior cases, the ellipsoidal truncation may miss significant posterior regions and include irrelevant, low-posterior-density areas. To overcome these difficulties, Metodiev et al. [12] subsequently adapted THAMES for multivariate mixture models, addressing key limitations of the earlier method, but the efficiency of their approach still depends on adaptation to the geometry of the posterior region and a potential computational cost arising from using auxiliary simulations to evaluate complex intersection volumes.
@@ 2 Approximating the evidence by harmonic means
and it holds for all probability density functions $\varphi(\cdot)$ that are defined over the posterior support. This identity justifies the following estimator
@@
Instrumental prior distribution
@@
Gaussian instrumental distribution
@@
$V(A)=\frac{\pi^{d/2}r^{d}|\hat{\Sigma}|^{\frac{1}{2}}}{\Gamma(d/2+1)},$ (4)
@@
Uniform instrumental distribution for a convex hull of $\alpha$-HPD samples
In order to bound the variance of the estimator (2), Robert and Wraith [21] proposed choosing the instrumental function $\varphi(\cdot)$ as a uniform distribution over a HPD region of coverage $\alpha$ (called an $\alpha$-HPD region). This region is approximated from the $\alpha$ fraction of a simulated posterior sample that corresponds to the largest arguments of $\pi(\theta)L(\theta)$ by constructing a convex hull, $\mathfrak{H}_{\alpha}$. The associated estimator (2) is then
@@
The authors illustrated the efficiency of the method on a $2$-dimensional toy example. However, the extension to higher dimensions is uncertain, due to the computational challenge of deriving the volume of the convex hull. In addition, the inclusion of the hull into an HPD region is not guaranteed and may be inefficient.
Partition Weighted Kernel (PWK) instrumental distribution
Wang et al. [22] introduced the partition weighted kernel (PWK) estimator, which generalizes the harmonic mean (HM) and inflated density ratio (IDR) methods via local weighting over a partition of the parameter space. Let $\Omega\subset\Theta$ be a working domain where the kernel $q(\theta)=\pi(\theta)L(\theta)$ is bounded away from zero, and let $\{A_{1},\dots,A_{K}\}$ partition $\Omega$ with weights $w_{k}>0$. The instrumental function is:
$\varphi(\theta)=\sum_{k=1}^{K}w_{k}\mathbf{1}\{\theta\in A_{k}\}.$
The PWK estimator is then:
$\hat{Z}^{-1}=\frac{\frac{1}{T}\sum_{t=1}^{T}\frac{\varphi(\theta^{(t)})}{q(\theta^{(t)})}}{\sum_{k=1}^{K}w_{k}V(A_{k})},$
where $V(A_{k})$ is the volume of $A_{k}$. Variance is minimized when $q(\theta)$ is nearly constant on each $A_{k}$, achieved in practice by standardizing the MCMC sample and using concentric spherical shells with representative kernel values at shell midpoints. The estimator is consistent and has finite variance under mild conditions. However, it becomes computationally intensive in high dimensions due to volume calculations and partition refinement, resembling a grid-based approximation.
To handle multimodal or skewed posteriors where constant weights over spherical shells may be inefficient, they proposed an extended version (ePWK) that allows functional weights $w_{k}(\theta)$ varying within each partition subset. In practice, this is achieved by further slicing each spherical shell $A_{k}$ into $m_{k}$ angular subsets $\{A_{k1},\dots,A_{km_{k}}\}$ and assigning a representative kernel value $q(\theta^{*}_{k\ell})$ to each slice. This additional partitioning reduces heterogeneity of the kernel within each subset, improving estimator efficiency for complex posterior geometries, though the computational overhead limits its practicality to lower-dimensional settings.
Uniform instrumental distribution over a Gaussian ellipsoid
Related to the proposal of DiCiccio et al. [3], Metodiev et al. [13] developed an instrumental distribution called Truncated Harmonic Mean Estimator (with the acronym THAMES). THAMES is a combination of both methods previously mentioned and uses a uniform distribution over the ellipsoid $A$ defined in (3). Since the volume of the ellipsoid $A$ is available as (4), THAMES bypasses both the computing issue of the volume and the derivation of the convex hull of Robert and Wraith [21]. Like the previous estimators, THAMES provides an unbiased estimator of $Z^{-1}$, provided that the posterior density does not vanish within the ellipsoid $A$. While the choice of the radius $r$ towards minimizing the variance of $\hat{Z}^{-1}$ is argued to be about $\sqrt{d+1}$ as in the Gaussian case, there is no guarantee that $A$ is included within a HPD region. Therefore, it faces limitations in multimodal cases (a case illustrated later on Figure 7), which may invalidate THAMES’ consistency and reliability in these settings.
Uniform instrumental distribution over a truncated Gaussian ellipsoid
Most recently, Metodiev et al. [12] proposed a new version of THAMES, which we denote by tTHAMES (truncated THAMES) for simplicity, and is specifically tailored for multimodal posterior distributions. This updated method relies on Reichl [18]’s proposal and includes a truncation strategy, towards using a support $A^{\prime}_{\hat{\theta},\hat{\Sigma},r,\alpha}$ made of the intersection of the ellipsoid $A$ in (3), and a $50\%$-HPD region. The THAMES for a mixture of $j=1,\ldots,J$ distributions of the same family,
@@
$A^{\prime}_{\hat{\theta},\hat{\Sigma},r,\alpha}=\{\theta:(\theta-\hat{\theta})^{T}\hat{\Sigma}^{-1}(\theta-\hat{\theta})<r^{2},\pi(\theta)L(\theta)>\hat{q}_{\alpha}\}$ (6)
@@ 3 ECMLE method
The ECMLE estimator is based on the Gelfand-Dey identity (1) and the framework of Robert and Wraith [21] (5), where a uniform distribution over a convex hull is used as the instrumental function. In contrast, our method constructs the set $\mathcal{E}$ as a union of disjoint, locally adapted ellipsoids, enabling exact volume computation while effectively capturing multimodal and irregularly shaped posteriors.
@@
Unbiased estimation requires that the region $\mathcal{E}$ is constructed independently of the sample used for evaluation. Given a collection of posterior draws $\{\boldsymbol{\theta}_{i}\}_{i=1}^{2T}$ from either MCMC or independent sampling, we partition them into two equal subsets. The first subset $\{\boldsymbol{\theta}_{t}\}_{t=1}^{T}$ is used to construct the disjoint ellipsoid covering of the HPD region, while the second $\{\boldsymbol{\theta}_{t}\}_{t=T+1}^{2T}$ evaluates the estimator.
Given the chosen HPD level $\alpha$, we determine a threshold $c$ on the unnormalized posterior densities $\tilde{\pi}(\theta|x)=\pi(\theta)\,L(x|\theta)$ that separates the high- and low-density samples (Algorithm 1). This yields the sets of high-density points $\boldsymbol{\Theta}_{\mathrm{HPD}}=\{\boldsymbol{\theta}_{t}:\tilde{\pi}(\theta_{t}|{x})\geq c,\ t\leq T\}$ and low-density points $\boldsymbol{\Theta}_{\mathrm{LPD}}$. The threshold $c$ serves as the target boundary that ellipsoids must respect.
@@
1:Input: HPD samples $\boldsymbol{\Theta}_{\text{HPD}}$ with posterior values, LPD samples $\boldsymbol{\Theta}_{LPD}$, posterior function $\pi(\boldsymbol{\theta})$, HPD level $\alpha$
2:Output: Union of non-overlapping ellipsoids, $\mathcal{E}$, and total volume $V(\mathcal{E})$
@@
6:Initialize: $\mathcal{E}\leftarrow\emptyset$, $V(\mathcal{E})\leftarrow 0$, $\mathrm{Centers}\leftarrow$ subsampled $\boldsymbol{\Theta}_{HPD}$
@@
27:  $V(\mathcal{E})\leftarrow V(\mathcal{E})+v$
@@
31:Return $\mathcal{E},\,V(\mathcal{E})$
We first subsample $\boldsymbol{\Theta}_{\mathrm{HPD}}$ by a rate $k$ (typically 0.05–0.1) to obtain candidate centers, ordered by posterior density to prioritize high-probability regions. For each candidate $\boldsymbol{\theta}^{*}$, we determine the ellipsoid shape through the following procedure:
A primary axis direction $\mathbf{u}_{1}$ points toward the nearest low-density point, capturing the dominant direction of posterior decay. Along this axis, we find radius $r_{1}$ where $\tilde{\pi}(\boldsymbol{\theta}^{*}+r_{1}\mathbf{u}_{1}|x)=c$ via a bisection search. An orthogonal basis $\{\mathbf{u}_{1},\ldots,\mathbf{u}_{d}\}$ is found using Gram-Schmidt, and semi-axes lengths $s_{i}$ are determined by finding the HPD boundary along each direction, yielding a shape matrix $\boldsymbol{\Sigma}=\mathbf{U}\,\mathrm{diag}(s_{1}^{2},\ldots,s_{d}^{2})\,\mathbf{U}^{\top}$. The resulting ellipsoid is then defined as
$\mathfrak{e}(\boldsymbol{\theta}^{*},\boldsymbol{\Sigma})=\{\boldsymbol{\theta}:(\boldsymbol{\theta}-\boldsymbol{\theta}^{*})^{\top}\boldsymbol{\Sigma}^{-1}(\boldsymbol{\theta}-\boldsymbol{\theta}^{*})\leq 1\}.$
To preserve the non-overlapping property required for volume computation, we adopt a two-step filtering procedure. Candidate centers are processed in order of decreasing posterior density. First, before constructing a new ellipsoid at candidate center $\boldsymbol{\theta}^{*}$ with maximum semi-axis $s_{\max}$, we check against all existing ellipsoids: if the Euclidean distance to any existing center $\boldsymbol{\mu}_{j}$ is less than the sum of maximum semi-axes (i.e., $\|\boldsymbol{\theta}^{*}-\boldsymbol{\mu}_{j}\|<s_{\max}+s_{\max,j}$), the candidate is rejected. Second, after an ellipsoid $\mathfrak{e}(\boldsymbol{\theta}^{*},\boldsymbol{\Sigma})$ is accepted, all remaining candidate centers that fall inside this ellipsoid are pruned from further consideration. This construction guarantees disjointness and enables the exact computation of the total volume as $V(\mathcal{E})=\sum_{j}\frac{\pi^{d/2}}{\Gamma(d/2+1)}\sqrt{\det(\boldsymbol{\Sigma}_{j})}.$
@@
With the ellipsoid collection $\mathcal{E}$ and its total volume $V(\mathcal{E})$ fixed, the marginal likelihood is evaluated using the second sample. For each parameter vector $\{\boldsymbol{\theta}_{t}\}_{t=T+1}^{2T}$, membership in $\mathcal{E}$ is determined by verifying whether it falls within ellipsoid $\mathfrak{e}_{j}$. The corresponding estimator is summarized in Algorithm 3.
@@
1:Input: Second sample $\{\boldsymbol{\theta}_{t}\}_{t=T+1}^{2T}$ with unnormalized posterior densities $\{\tilde{\pi}(\theta_{t}|{x})\}_{t=T+1}^{2T}$, ellipsoid collection $\mathcal{E}$, total volume $V(\mathcal{E})$
@@
$\hat{Z}^{-1}=\frac{1}{T}\sum_{t=T+1}^{2T}\frac{\mathbf{1}_{\mathcal{E}}(\theta_{t})}{V(\mathcal{E})\;\tilde{\pi}(\theta_{t}\mid x)}.$
@@
The computational efficiency of ECMLE is critical for practical applications. We analyze here the time complexity of each algorithm component, demonstrating that the method scales favorably with sample size and dimension.
@@
4 Numerical Illustrations
In this section, we empirically compare ECMLE with THAMES estimators through several examples. We focus on THAMES because it represents the current state of the art among harmonic-mean–based estimators. THAMES and tTHAMES were recently shown to outperform earlier bounded harmonic mean approaches and to provide stable, closed-form evidence approximations across a range of models. As ECMLE builds directly on the same harmonic-mean identity and HPD-based truncation principle, this comparison allows a fair and direct assessment within the same methodological family.
We use the notation $\boldsymbol{x}=\{x_{1},\ldots,x_{n}\}$ to indicate a set of $n$ observations. Each experiment was performed using 100 replications of Algorithms 1–3 for all examples. To ensure a fair comparison across methods, we calibrated the number of posterior draws for each method so that all methods operate under approximately the same computational budget. This approach accounts for the differing per-sample computational costs of each estimator, allowing us to compare estimation accuracy at matched runtime rather than matched sample size.
Example 1: Multivariate Gaussian distributions